OpenStudy (anonymous):

A rational function g is given below. Find all values of a for which g(a)=12. g(x)= 4/x+3 + 16/x^2+6x+9

OpenStudy (anonymous):

\[4/a+3+16/a^2+6a+9=12\\ \frac{4a+16}{a^2}=0\\ a=4 \]

OpenStudy (anonymous):

4/(a+3) + 16/(a+3)^2 = 12 4[1/(a+3) + 4/(a+3)^2] = 12 now taking lcm (a+3)+4= 3(a+3)^2 simplifying 3a^2 +17a +20 = 0 now using the quadratic formula to find the roots we will get a = -4 or -5/3

OpenStudy (anonymous):

thats what the question says?

OpenStudy (anonymous):

yea